OpenStudy (anonymous):

the difference between two numbers is 4. three times the smaller number minus four times the larger is -28.. what are the numbers?

OpenStudy (anonymous):

x = 16 y = 12

OpenStudy (anonymous):

9a squared - 121

OpenStudy (anonymous):

x - first number, y - second number \[\begin{cases}x-y=4\\3y-4x=-28\end{cases}\] I hope you can do it yourself now x=16 y=12

OpenStudy (aroub):

The first number: x ==> smaller number The second number: x+4 ==> larger number =>3(x)-4(x+4)=-28 =>3x-4x-16=-28 =>3x-4x=16-28 =>-x=12 =>x=12 The number are: 12,16 :D

OpenStudy (aroub):

numbers*

OpenStudy (anonymous):

i feel i had told u ld u the answer

OpenStudy (anonymous):

12,16 are not possible answers and i typed the qestion exactly as it reads. answers are, 5,9 8,12 7,11 6,10

OpenStudy (aroub):

Not possible!

OpenStudy (anonymous):

the problem you typed is definitely 12 and 16, maybe you typed it wrong?

OpenStudy (anonymous):

nope, read it word for word

OpenStudy (anonymous):

well, try plugging 12 and 16 into the problem you typed and you'll see they work correctly

OpenStudy (anonymous):

where from u got this question

OpenStudy (anonymous):

they do work

OpenStudy (anonymous):

i think answers are wrong

OpenStudy (aroub):

Yes, maybe